OpenStudy (mindblast3r):

A children's cough medicine requires 5mL of medicine for children 2 to 5 years old, and 7.5 mL of medicine for children between 5 and 12 years of age. Is this situation non-proportional or proportional?

OpenStudy (mindblast3r):

proportional as in like are they equal?

OpenStudy (rational):

check if below are true : \[\dfrac{2~years}{5~mL} = \dfrac{3 ~years}{5~mL} = \dfrac{6~years}{7.5~mL}\]

OpenStudy (rational):

If all the ratios above are same, then the medicine is proportional to age. otherwise its not proportional

OpenStudy (mindblast3r):

ohh, i get it now thank you very much :)

OpenStudy (rational):

yw